### v1

# Aragog: Just-in-Time Model Routing for Scalable Serving of Agentic Workflows

Abstract
Agentic workflows have emerged as a powerful paradigm for solving complex, multi-stage tasks, but serving them at scale is computationally expensive given the many LLM inferences that each request must pass through. Configuration selection, or the cost-aware assignment of workflow agents to specific LLMs, can reduce these costs, but existing approaches bind configuration decisions before request execution, making them ill-suited for the heterogeneous and lengthy execution of workflows. Specifically, system loads can fluctuate rapidly and substantially during a request’s lifetime, causing fixed configurations to quickly become suboptimal. We present Aragog, a system that progressively adapts a request’s configuration throughout its execution to match runtime dynamics. To make this practical despite the massive space of workflow configurations, Aragog decouples the problem into two core elements – a one-time routing step that identifies all accuracy-preserving configurations, and a cheap per-stage scheduler that selects among them using up-to-date system observations – and introduces novel strategies to accelerate each. Across diverse workflows and model families, Aragog increases maximum serving throughput by 50.0–217.0% and reduces median latency by 32.5–78.9% at peak request rates, while maintaining accuracy comparable to the most expensive configurations.
Aragog: Just-in-Time Model Routing for Scalable Serving of Agentic Workflows
Yinwei Dai§   Zhuofu Chen§   Anand Iyer   Ravi Netravali§
§Princeton University   Georgia Institute of Technology
1 Introduction
Autonomous agents, or Large Language Models (LLMs) augmented with tools, memory, and planning, have shown impressive ability to tackle complex tasks autonomously [yang2024sweagent, snowflake-agentic-workflow]. However, fully agentic systems often lack the structure and reliability needed in production [multiagentllmsystemsfail]. This has led to the rise of agentic workflows: directed graphs that orchestrate multiple agents [langchain, dspy, autogen, metagpt], each responsible for a well-defined subtask and passing outputs downstream. These workflows preserve the generative strengths of agents while offering the observability and debuggability of traditional software. Yet serving them at scale remains expensive, largely due to the inflated number of LLM invocations per request. A single workflow execution typically triggers several agent invocations, each making multiple LLM calls; the net effect is that LLM inference can account for nearly 80% of end-to-end request time in workflows (§2.1).
The predominant strategy for reducing these inference costs focuses on intelligently selecting workflow configurations – i.e., end to end assignments of specific LLMs to each agent in the workflow – that minimize cost while preserving accuracy. Existing methods operate at one of two granularities: (1) per-workflow optimizations [cognify, llmselector, murakkab] use historical data to select configurations offline for all inputs, while (2) per-input optimizations [masrouter, routellm, daao] use online routers to select configurations dynamically for each request. Both, however, optimize for static cost metrics (e.g., API dollars or FLOPs) and bind configurations before processing a request, ignoring the increased dynamism that agentic workflows bring to serving systems. Specifically, serving a request is both heterogeneous and lengthy. As it moves through a multi-step workflow graph, the computational cost it brings varies across agents, and the system resources available to it rapidly fluctuate as other requests arrive and execute. Consequently, a configuration selected at the start may quickly become suboptimal during a request’s lifecycle, e.g., an agent assigned a small model to save FLOPs may perform worse than necessary (latency- and throughput-wise) if that model is overloaded and a larger model is idle when the request reaches the agent. Across agentic workloads (§2.2), we find that such static strategies can miss 42.8–233.3% of the potential gains from configuration selection.
To realize the promise of configuration selection, our guiding observation is that the same multi-stage structure that increases the cost of agentic workflows also creates unique opportunities for runtime optimization that existing systems overlook. Workflows exhibit increased fault tolerance – errors made early can often be corrected downstream, e.g., via refinement agents – which broadens the set of viable configurations for each input (configuration flexibility). Further, their stage-by-stage [^1] In this paper, we use the terms stage and agent interchangeably to refer to individual execution steps in an agentic workflow. execution also enables configurations to be adapted as system states evolve, rather than fixed upfront (scheduling flexibility). However, exploiting these flexibilities is challenging. Identifying accurate configurations requires exploring a combinatorial space that grows exponentially with workflow depth and model choices, making per-stage exploration prohibitively expensive. Moreover, stage-wise model decisions must be coordinated across concurrent inputs to account for shared resource contention, further compounding the complexity as more requests and configuration options are introduced.
We present Aragog, a workflow serving system that overcomes these challenges by decomposing configuration selection into two complementary steps. Our key insight is that, for a given input, the accuracy of a configuration is fixed, whereas its runtime performance can change dramatically as system load fluctuates. Aragog leverages this asymmetry by performing the expensive, accuracy-related analysis once per request (i.e., when it arrives) to identify all configurations that preserve accuracy. Then, as the request moves through the workflow, Aragog rapidly selects among only the pre-validated configurations for each upcoming stage, using current runtime signals to stay aligned with system conditions.
To realize this decoupling approach while maximizing serving throughput and avoiding any added latency from configuration decisions, Aragog embeds two optimized components. First, to identify the set of accuracy-preserving configurations, Aragog decomposes the routing problem into many simple binary classification tests and exploits the near-monotonic relationship between configuration FLOPs and accuracy to prune most candidates quickly; this keeps routing both accurate and lightweight enough to overlap with request queuing delays. Second, Aragog’s (stage-wise) runtime scheduler leverages how FIFO priorities and pruned configuration sets naturally structure the assignment space: earlier requests in the queue have only a few viable configurations – since each choice at an earlier stage can invalidate options that use a different model for that stage – and once chosen, they constrain later requests rather than letting the space fan out. This, in turn, creates a compact search space that beam search can traverse efficiently, avoiding both exhaustive exploration and greedy myopia.
We evaluate Aragog by comparing with existing configuration optimizers – both per-workflow [cognify, llmselector, murakkab] and per-input [routellm, masrouter, daao] – augmented with oracle accuracy information about configurations, i.e., representing their best possible performance. Across five popular agentic workflows and multiple model families (Qwen 2.5 [qwen], Llama 3 [llama3], and Phi 4 [phi4]), we find that Aragog improves maximum serving throughput by 50.0–76.3% over per-input optimizations and 78.1–217.0% over per-workflow optimizations. Further, Aragog consistently reduces latency across varying request rates. For instance, at peak load, median and P95 reductions were 32.5–58.8% and 46.8–66.1% over per-input optimizations, and 60.0–78.9% and 63.2–81.5% over per-workflow optimizations. Crucially, Aragog achieves these performance wins while maintaining accuracy within 1% of that when always using the most expensive configuration.
2 Background and Motivation
We begin by providing an overview of agentic workflows (§2.1), highlighting the limitations of current optimization approaches (§2.2). We then explore the promising opportunities from capitalizing on the unique runtime flexibility of agentic workflows and discuss the practical challenges in effectively realizing these opportunities.
2.1 Agentic Workflows
Agentic workflows are directed graphs of agent invocations, where each agent—a prompted LLM augmented with tools, memory, and planning capabilities—executes a well-scoped sub-task and passes its output to downstream agents. This structured approach offers similar observability and debuggability expected from conventional software while leveraging LLMs’ generative capabilities—advantages that fully autonomous agents cannot provide [multiagentllmsystemsfail]. As shown in Figure 1, example patterns and application domains include: (1) self-refinement workflows [self_refine] that iteratively improve outputs through generation, critique, and refinement agent; (2) natural language to SQL systems [chess, hexgentext2sql] that decompose query construction into keyword extraction, column selection, SQL generation, and refinement stages; (3) task decomposition workflows [decompose] that split complex problems into parallel sub-tasks handled by different agents before aggregating results; and (4) multi-agent voting systems [self-consistency, voting] that ensemble multiple agents for robust decision-making. In practice, these workflows are explicitly encoded using frameworks like LangChain [langchain] and DSPy [dspy]. Once a workflow graph is defined, developers can specify a configuration that assigns each agent to a chosen LLM, enabling accuracy-cost tradeoffs without modifying workflow logic. These configurations often utilize heterogeneous models [llmselector]–for example, using smaller models for easier agent tasks and larger models for the rest—with each model hosted on different serving engines across different GPUs.
Yet despite these intuitive benefits, agentic workflows face significant practical challenges, largely centered around the high costs they bring. Indeed, each agent invocation incurs the inference cost of powerful yet resource-intensive LLM calls. These expenses scale quickly since workflows generally involve multiple sequential or parallel agent calls. For instance, coding workflows require 3–5 agent calls per query to plan, generate, test, and verify solutions [zhang2023agents, yang2024sweagent], while even simple workflows such as multi-hop question answering for legal or scientific search can involve 5+ agent calls [lewis2020rag, bai2023constitutional, orca]. Across our workloads, LLM calls account for 79.4% of end-to-end latency on average, with tool calls and framework overhead accounting for the remaining 20.6%.
Figure 1: Examples of common agentic workflows that vary in terms of both structure and application domains.
2.2 Limitations of Existing Optimizations
Existing systems focus on configuration selection to reduce the core LLM inference costs while preserving accuracy, operating at per-workflow or per-input granularity. We categorize and discuss these techniques, highlighting their limitations; complementary optimizations are covered in (§6).
Per-workflow optimization. Systems like Cognify [cognify], LLMSelector [llmselector], and Murakkab [murakkab] perform one-time or periodic configuration selection based on sample data, minimizing costs (e.g., FLOPs, API costs) while preserving accuracy relative to the most expensive configuration. Once selected, a configuration is fixed for all requests, eliminating runtime overhead but sacrificing per-input adaptability.
Per-input optimization. In single-LLM serving, existing LLM routing approaches [routellm, mixllm] leverage input heterogeneity by routing it to the most cost-efficient model that meets accuracy requirements. Recent works extend this routing paradigm to agentic workflows [daao, masrouter] by selecting not just singular models, but a workflow configuration for each input. Routers (e.g., transformer-based) are trained using accuracy labels relative to the most expensive configuration, enabling cost reduction while preserving accuracy.
The problem. Though effective, existing optimizations–targeting per-workflow or per-input granularity fail to adapt to the runtime dynamics (e.g., resource availability). The core issue is that they bind configurations before execution and select configurations based on static cost metrics (e.g., FLOPs). As a result, configurations become stale by execution. Worse, the long and multi-stage nature of agentic workflows exacerbates this staleness, as system states can shift dramatically throughout their execution.
Binding too early to configuration selections. Existing approaches all make static configuration decisions prior to any workflow execution, i.e., either once before all inputs, or at the start of each input. The core issue is that system states change rapidly, rendering these a priori selections stale by execution time. Natural LLM dynamics already introduce significant variance: output lengths vary unpredictably, queue depths fluctuate continuously, and dynamic batching reshapes execution efficiency moment by moment [orca]. Agentic workflows amplify this unpredictability by compounding variance across multiple LLM calls. Consequently, any upfront configuration selection before execution can quickly become stale by the time it executes.
Optimizing for static metrics. Current systems minimize predetermined costs like FLOPs or API call prices, but these fail to translate to actual runtime costs, i.e., on metrics such as resource utilization and latency. Instead, what truly governs these aspects is how configurations fit into the current system states and utilize the currently available resources. For instance, a “more expensive” 70B-parameter model on an idle GPU can return results faster than a “cheaper” 14B model with a saturated request queue.
Figure 2: Static approaches (per-workflow, per-input) achieve suboptimal maximal serving capacity. Incorporating runtime cost metrics and runtime configuration adaptation incrementally improves serving capacity. Results are evaluated on HumanEval [humaneval], Bird [bird], and StrategyQA [strategyQA].
To quantify the inefficiency, we follow the experiment setup from (§5.1) and evaluate three workflows–self-refine, text-to-sql, and task decomposition–with different datasets in Figure 1 using three model candidates: Qwen {7,14,32}B [qwen]. For space, we primarily use these workflows here, but note that the trend holds for other workflows and workloads used in our evaluation. In Figure 2, we compare serving capacity (maximum throughput, normalized to the best approach) across configuration selection strategies on three workloads. All approaches are given perfect accuracy knowledge to isolate their fundamental limitations. Per-workflow optimization achieves only 30–45% of maximum capacity. Per-input optimization, which selects the lowest FLOPs accurate configuration for each input, improves serving capacity to 54–63%, demonstrating the benefit of per-input adaptability. Per-input++ is further augmented with a runtime cost oracle that incorporates runtime system metrics into its cost estimation. This improves serving capacity to 70-75%, demonstrating the benefits over using static cost metrics. Per-input++++ further adds a runtime configuration adaptation oracle, adapting configuration based on fresh system states rather than fixing configurations throughout execution. This achieves the highest serving capacity, leaving a 25–30% gap over Per-input++. This significant gap highlights the benefits of continuous and rapid configuration adaptation at runtime over static, upfront configuration selection.
Despite inflating configuration staleness for existing optimizations, the multi-stage structure of agentic workflows also provides unique opportunities to achieve runtime configuration adaptation. Concretely, we identify two forms of flexibility inherent to agentic workflows that enable adaptation to rapid system dynamics.
Configuration Flexibility: Despite the additional complexity and overhead it brings to configuration selection, the exponentially large search space for agentic workflows also presents remarkable error tolerance–when upstream agents produce suboptimal outputs, downstream agents can recover from these intermediate errors (Figure 3). This self-correcting behavior dramatically expands the viable configuration space: for each input, numerous configurations can achieve the target accuracy, not just one “optimal” choice (Figure 4). This large candidate space makes it possible to cater to rapid system dynamics and achieve per-input configuration selection that aligns with dynamic runtime system state, bypassing the limitations of existing routers trained with static cost and upfront fixed configurations.
Scheduling Flexibility: The multi-stage structure of workflows enables dynamic configuration switching during execution. Unlike traditional single LLM systems that have to commit to full configurations upfront, the multi-stage structure naturally allows configuration selection to evolve as the workflow progresses. Specifically, this stage-by-stage execution makes it possible for runtime reconfiguration, which can switch between configurations that share the same prefix–completed stages remain valid while downstream agents can be reassigned based on current system states.
These two forms of flexibility–multiple viable configurations per input and progressive scheduling–highlight the untapped potential in existing systems to adapt configuration to runtime dynamics. However, exploiting this flexibility requires solving fundamental systems challenges in managing the runtime overheads, given the large configuration space and the need for frequent runtime reconfiguration.
Figure 3: Agentic workflows exhibit robustness by recovering from intermediate errors. For each workflow configuration, we track the inputs with intermediate errors, and record the fraction subsequently corrected by the remained workflow. The plot shows the CDF of these correction percentages across all configurations.
Figure 4: Numerous workflow configurations can yield accurate results at per-input granularity.
[…]
C2: Joint runtime reconfiguration under concurrency. Beyond per-input routing overheads, runtime reconfiguration amplifies the coordination that serving platforms must consider when managing concurrent requests. Specifically, capitalizing on each request’s routing flexibility can constrain other requests. For instance, consider two requests where R1 can achieve target accuracy with either 7B or 14B models, while R2 needs either 7B or 32B. If the serving engine for the 7B model can only house one more request, independent scheduling might grant R1 use of the 7B model first (its cheapest option), forcing R2 to use the expensive 32B model. Joint scheduling would recognize that R1 can make a slight performance sacrifice by using 14B, allowing R2 to take the scarce 7B slot—improving overall system throughput and latency. However, such joint optimization becomes computationally expensive as it scales exponentially with the number of concurrent requests and configurations.
3 Design
We present Aragog, an efficient serving system for agentic workflows that dynamically selects workflow configurations and progressively schedules their execution based on real-time system state. Its overarching goal is to optimize the serving capacity of agentic workflows while preserving accuracy. To realize that, the key insight is that the accuracy of a configuration for a given input remains constant, while its runtime cost rapidly fluctuates with system dynamics (i.e., resource availability). Traditional monolithic routers jointly evaluate both accuracy and cost, necessitating frequent and expensive router invocations to account for runtime cost. By recognizing this fundamental asymmetry between accuracy (static and expensive) and cost (dynamic and cheap), we decouple them to make frequent runtime reconfiguration practical: the expensive accuracy evaluation is performed once upfront to identify the viable configuration set, while the lightweight cost-based selection happens frequently at execution time when actual costs are observable. This enables tractable per-input configuration selection (C1) and joint optimization across concurrent requests (C2) at scale.
Specifically, Aragog decouples configuration selection into two components: (1) A configuration predictor3.1) that identifies all accurate configurations for each input before execution, rather than a single configuration. (2) A runtime scheduler3.2) progressively selects from these candidates stage-by-stage at execution time, efficiently coordinating configuration selection across concurrent requests to maximize serving capacity while preserving each request’s flexibility. Figure 5 overviews Aragog’s workflow. 1 Follow the existing systems, Aragog trains configuration predictors offline using representative data. 2 At runtime, incoming queries pass through the configuration predictor to obtain their viable configuration set. 3 Meanwhile, the runtime scheduler monitors the available resources in each model’s serving engine. 4 With this runtime information, the scheduler performs stage-wise scheduling, jointly selecting configurations across all concurrent requests from their respective configuration sets based on slot availability as each request progresses through stages. 5 After each stage, Aragog prunes configurations with different prefixes, progressively narrowing viable options.
Figure 5: Aragog predicts a set of accurate configurations for each input before execution and performs stage-wise scheduling at runtime. At stage 0 example, model C is overloaded, forcing query 2 to use model A. To accommodate this, the scheduler exploits query 1’s flexibility, picking configurations that start with model B.
3.1 Efficiently Predicting Accurate Configuration Sets
To capitalize on the runtime flexibility of agentic workflows, Aragog’s configuration predictor fundamentally differs from traditional routers–instead of selecting a single best configuration, it must identify the set of accurate configurations for a given input $x$. This configuration set, denoted $\mathcal{C}(x)$, is essential for runtime scheduling flexibility: the prediction must be cheap enough to avoid becoming a bottleneck, while remaining accurate and presenting the maximum number of configurations to maximize runtime flexibility.
Problem formulation. Given an input $x$ and a workflow DAG $G$ with $n$ agents, let $\mathcal{M}=\{m_{1},...,m_{k}\}$ denote the set of candidate models ordered by size (e.g., Qwen-7B through Qwen-32B), where $\mathcal{C}_{\mathrm{all}}=\mathcal{M}^{n}$ represents the full configuration space. We define $c^{*}$ as the most expensive configuration where all agents use the largest model $m_{k}$. Our objective is to find as many configurations as possible whose accuracy matches $c^{*}$ for input $x$, while minimizing overhead.
Decomposing into binary decisions. Existing routing approaches use a standard two-layer architecture: an embedding model (often transformer-based) encodes the input into a high-dimensional representation, followed by a classifier that predicts the configurations to select from $|\mathcal{M}|^{n}$ options. To practically deploy routers, existing systems navigate the tradeoff between router complexity and accuracy. However, identifying the set of viable configurations introduces a fundamentally different problem that opens a larger design space beyond single-router complexity tradeoffs. The design space spans a spectrum of granularities based on how we partition the configuration space. At one extreme, a single router evaluates all $|\mathcal{M}|^{n}$ configurations simultaneously; at the other, $|\mathcal{M}|^{n}$ binary routers each evaluate one configuration independently. Between these extremes lie intermediate approaches, such as $k$ routers each handling $|\mathcal{M}|^{n}/k$ configurations. The fundamental tradeoff is that larger subsets reduce router count but increase per-router complexity, making it difficult to reason about total routing costs.
To systematically explore this tradeoff, we evaluate how the overall routing accuracy and cost accuracy scale with subset size under a fixed parameter budget. We experiment with three workflows, and follow the router architecture in (§5.1). We partition these 16 configurations into different subset sizes: 16 routers handling 1 configuration each (binary), 4 routers handling 4 configurations each, and 1 router handling 16 configurations. As the subset size increases, we scale router parameters linearly to maintain constant total router parameters. Figure 6 shows the results averaged across inputs: Despite allocating the same total parameters across all approaches, the overall accuracy degrades with increasing subset size. Binary routers excel in both precision (fewer inaccurate configurations) and recall (more viable configurations for runtime scheduling). A collection of binary routers achieves 95–97% precision and 47-67% recall, while a single 16-configuration router with the same total parameters achieves only 21–40% precision and 11-21% recall.
This accuracy degradation occurs because routers must learn increasingly complex decision boundaries to distinguish multiple configurations simultaneously, which makes the collection of binary routers the most accurate with the same parameter budget. Therefore, Aragog adopts the finest granularity: decomposing the routing problem into a collection of binary routers $\{f_{c}\}_{c\in\mathcal{C}_{\text{pool}}}$, where each router $f_{c}$ makes a simple binary decision–is configuration $c$ as accurate as $c^{*}$ for input $x$.
Figure 6: Overall routing accuracy drops as each router’s configuration space increases. Binary routers achieve the best precision and recall for a given parameter budget. Our evaluation (§5) confirms that the recall is sufficient to provide flexible runtime scheduling.
With the offline trained binary routers, our goal is for routing to complete during queuing for the first stage of a given request. This requires router inference to be deadline aware–finishing within the request’s estimated queuing delay, as any computation beyond this threshold directly adds latency to that request’s end-to-end serving latency. However, naively running all routers for every possible configuration would itself exceed typical queuing delays. Fortunately, we observe that significant redundancy exists in the configuration space, which we exploit as a knob to dynamically adjust routing inference budget based on estimated queuing delays.
Pruning the configuration space. To reduce the router’s search space, we leverage a key empirical property of model capabilities. We observe that the accuracy of a workflow exhibits near-monotonicity under model upgrades: replacing any agent’s model with a more capable one (i.e., one with more parameters within the same model family) maintains or improves end-to-end accuracy in the vast majority of cases. Figure 7 validates this property empirically across all configurations of a self-refine workflow for the coding task, where green arrows indicate accuracy-preserving upgrades and red arrows indicate rare violations (only 8 out of 288 upgrade cases, with accuracy drops capped at 1.5%). We observe a similar trend consistently across different model families and agentic workflows in our evaluation. The intuition is straightforward: if we run a router on a configuration and it deems that configuration accurate enough, then any configuration that upgrades one or more agents to more capable models are expected to be accurate. Conversely, if a configuration is deemed inaccurate, downgrading any agent will likely remain inaccurate. Built on that, Aragog can effectively prune the whole configuration search space. Crucially, Aragog leverages this empirically observed (though non-guaranteed) property solely to prune the set of configurations; all selected configurations still run through the corresponding binary router before being considered during serving. Thus, violations of this property can only result in reduced serving flexibility by missing viable configurations, as inaccurate configurations will be filtered.
Concretely, the near-monotonic property allows us to avoid exhaustive router evaluation through a two-phase approach: targeted binary search followed by verification. We first construct the search space as a collection of monotonic configuration chains. Each chain starts from the base configuration (all agents using the cheapest model $m_{1}$). We construct these chains through depth-first search (DFS) by individually upgrading each agent’s model while keeping others fixed. Along each chain, we locate the transition boundary between inaccurate and accurate configurations via binary search. The boundary implies that subsequent upgrades are accurate (and all prior downgrades are inaccurate). We add the accurate configurations to our candidate set. This systematic search identifies all boundary configurations without exhaustive router evaluations, dramatically reducing our search from $|\mathcal{M}|^{n}$ total configurations to a much smaller candidate set. Again, before passing these candidates to the scheduler, we run their corresponding binary routers to confirm viability and filter out any false positives.
Figure 7: Workflow accuracy is largely monotonic under model upgrades. Green: valid monotonic upgrades; Red: rare violations (8 out of 288 model upgrades, with mild accuracy drops << 1.5%). The trend holds for other workflows.
Efficient router inference Even after pruning, naively evaluating remaining binary routers sequentially is inefficient. As described before, our router architecture follows the standard design used in existing systems: an embedding model (generates a general high-dimensional representation for each input) followed by task-specific classifiers (classify whether specific configurations are accurate). Though the general embedding model can be shared to maximize GPU utilization through batching. The problem is that the sequential execution of specialized classifiers wastes GPU resources–each head’s computational footprint cannot fully utilize GPU resources. To address this, we fuse classification heads into a single batched operation layer by layer through customized GPU kernels. This transforms $k$ sequential memory-bound operations into one compute-bound operation that fully utilizes the GPU compute and reduces kernel launch overheads.
Figure 8: Reduction in average per-input routing latency from cumulative optimizations.
Optimization effectiveness. Figure 8 shows the cumulative impact of our optimizations on average per-input routing latency for two representative workflows. Starting from naive exhaustive evaluation of $|\mathcal{M}|^{n}$ routers (baseline), the monotonic property reduces routing overhead by 30.0–47.6% by configuration pruning. Kernel fusion for classification heads further reduces latency by 33.1–56.3%. Together, these optimizations reduce average per-input routing latency from 90.9–196.2 ms to 42.6–44.9 ms per request.
Putting it all together: deadline-aware configuration search. Aragog adapts its search strategy to the expected budget–the time before the request’s first stage. The scheduler maintains exponential moving averages of per-request queuing times and sets the search deadline accordingly. When queues are deep (high load), longer wait times allow comprehensive search: we perform the full monotonic exploration to identify boundary configurations, then evaluate all remaining candidates with the corresponding binary classification routers. When queues are shallow (low load), we employ early stopping and terminate the search even if more configurations remain unexplored.
3.2 Runtime Configuration Selection and Scheduling
Given the set of viable configurations from §3.1 for each request and a set of model serving engines (with provisioned resources for each model that Aragog takes as input), Aragog’s goal is to leverage the configuration flexibility and make fine-grained, frequent configuration decisions to best adapt to system dynamics. To realize this, Aragog continuously monitors the available slots for each model’s serving engine. With this fresh view of system state and each request’s viable configurations, Aragog employs stage-wise scheduling that defers configuration selection until necessary, ensuring that selected configurations optimize for the actual resource conditions they encounter to minimize staleness. Moreover, Aragog jointly optimizes across multiple requests to mitigate the contention between concurrent requests, while bounding the coordination overheads (C2).
More specifically, with the following inputs: (i) $\mathcal{Q}=\{r_{1},r_{2},...,r_{n}\}$, the request queue in FIFO order where each $r_{i}$ is at a specific workflow stage; (ii) $\mathcal{C}(r_{i})$, the set of viable configurations for request $r_{i}$; and (iii) $S_{m}(t)$, the available slots for model $m$’s engine at time $t$ (considering current batch occupancy and maximum batch size). The expected output is an assignment of requests to models for the current scheduling round, with the objective of minimizing end-to-end latency and maximizing throughput while preserving the common FIFO ordering. Since each engine’s maximum batch size represents its saturation point, this objective reduces to maximizing utilization across all serving engines–undersaturated engines waste resources and increase queuing delays for pending requests.
Joint optimization across requests and time. Given the mentioned inputs, Aragog must assign each request in the FIFO queue to a configuration from its viable set. The quality of this assignment is critical: it not only leads to resource underutilization in the current round but also future scheduling flexibility, as each configuration selection prunes the configuration set. However, the assignment space is prohibitively large–exponential to the number of requests and the size of their configuration set–making exhaustive search impractical for real-time scheduling.
Our key insight is that FIFO ordering and shrinking configuration sets together narrow the assignment space for efficient exploration. First, FIFO ordering prioritizes earlier requests, and their assignments constrain the available options for later requests due to shared resources. Second, earlier requests in the queue are typically at later workflow stages with already-pruned configuration sets, leaving them with few viable options. Together, these two factors create a narrow search space: earlier requests have limited options, and once assigned, they further restrict later requests. This makes beam search ideal–it avoids both the intractability of exhaustive search and the myopia of greedy assignment by iteratively extending a small set of promising partial assignments in FIFO order.
Figure 9: Beam search balances quality and efficiency, achieving near-optimal assignments comparable to brute-force while maintaining efficiency close to greedy search. Bars list averages across workloads, with error bars for min-max.
Aragog maintains only the top-$B$ (B is defaulted to 4 and evaluate different values in §5.3) most promising partial assignments at each step, where each partial assignment represents a subset of scheduled requests with their configuration assignments. Partial assignments are ranked by two criteria: (1) weighted utilization—filling model slots proportional to each engine’s profiled throughput (e.g., a 32B model slot contributes less than a 7B slot with same amount of serving resource)—and (2) flexibility preservation as a tiebreaker, the percentage of configuration options preserved for request’s future stages. This balances immediate resource contention with preserving a request’s future flexibility.
At each step, the scheduler expands partial assignments in FIFO order by considering all valid configuration assignments for the next unscheduled request. To avoid blocking, we employ look-ahead FIFO: when a request can not be scheduled due to resource constraints, the scheduler skips it and proceeds to the next request in the queue. This is critical for heterogeneous workflows where requests require different models–strict FIFO would block subsequent requests with flexible configurations behind a constrained request waiting for an overloaded model. Our relaxation preserves FIFO fairness because: (1) skipped requests add no additional delay–they remain blocked regardless of whether subsequent requests are scheduled, and (2) Aragog still iterates through requests in FIFO order in future rounds, ensuring skipped requests are never overtaken by later joint requests requiring the same models.
By retaining only $B$ states throughout the search, Aragog makes joint optimization tractable while finding high-quality assignments. The search terminates when all feasible requests are assigned or no valid extensions exist. To evaluate beam search’s performance, we take periodic snapshots of our evaluation workloads during serving. Figure 9 shows that beam search achieves near-optimal assignment quality–measured by resource utilization and flexibility preservation–comparable to exhaustive search while maintaining efficiency close to greedy approaches (order of magnitude faster than brute-force).
Supporting complex workflows. Agentic workflows extend beyond sequential workflows, often containing parallel branches and complex dependencies. Aragog leverages topological sorting to ensure correct execution order while maximizing parallelism. Before execution, Aragog performs a topological sort on the workflow DAG and pre-computes each agent’s depth—the longest path from that agent to any leaf node. During execution, similar to existing systems [parrot, ayo], Aragog follows a two-level priority scheme: (1) inter-request: FIFO ordering prioritizes earlier requests over later ones, and (2) intra-request: agents are prioritized by depth in descending order to schedule critical-path agents first. For example, given two requests $R_{1}$ (arrived first) and $R_{2}$ (arrived second), where $R_{1}$ has ready agents $A$ (depth=4) and $B$ (depth=2), and $R_{2}$ has ready agent $C$ (depth=5), the scheduling order is: $A$ from $R_{1}$ (highest depth in earlier request), then $B$ from $R_{1}$ (same request), and finally $C$ from $R_{2}$ (later request), despite $C$ having the globally highest depth. With this ordering, Aragog’s optimization applies seamlessly to complex workflows.
4 Implementation
class Workflow(ABC):
@abstractmethod
def extract_workflow_graph(
self, dspy_program: dspy.Module
"""Extract workflow DAG from DSPy program."""
pass
@abstractmethod
async def execute_stage(
self,
stage: int,
model_idx: int,
lm: dspy.LM,
request_state: RequestState
"""Execute a stage with the specified model."""
pass
Figure 10: Aragog’s DSPy program wrapper abstraction for enabling graph structure extraction and stage-wise scheduling.
We implement Aragog as an orchestration layer in 6.8K lines of Python and Triton code that sits between agentic workflow front end and LLM serving engines without the need of modifying both.
Frontend and Workflow Specification. As shown in Figure 10, Aragog converts any DSPy program into its internal workflow representation, enabling runtime model reconfiguration and asynchronous stage-wise execution. We choose DSPy for its lightweight design and ease of use, making it popular for building agentic workflows. However. Aragog can support other agentic frameworks with a specific implementation to extract the workflow graph.
Router Inference We implement dynamic batching for shared embedding model serving. To accelerate classifier inference, we develop custom Triton kernels that fuse classifiers across a tunable number of classifiers into a single batched operation. We leverage CUDA graphs to further reduce the kernel launching overheads.
Runtime Scheduler. The scheduler runs asynchronously. It maintains: (1) FIFO request queues with look-ahead scheduling, (2) per-model engine states tracking batch occupancy, (3) configuration caches mapping requests to viable sets, and (4) exponential moving averages of queue metrics. We implement beam search using priority queues, where each state encodes partial assignments as bit vectors for efficient manipulation.
Serving Backend Integration. Aragog use SGLang [sglang], a widely used serving framework for large language models as our serving backend. However, Aragog naturally supports any serving backends that provide OpenAI-compatible APIs (vLLM [vllm], TensorRT-LLM [tensorrt-llm] and HuggingFace TGI [TGI]).
5 Evaluation
We evaluated Aragog across a wide range of agentic workloads and varying request rates. Our key findings are:
  • $\bullet$
    Aragog improves maximum serving throughput by 50.0–76.3% and 78.1–217.0% over per-input and per-workflow optimizations, respectively.
  • $\bullet$
    Aragog maintains lower latency under increasing load. Under the highest load for all workloads, Aragog reduces P25, median, and P95 end-to-end latency by 23.6–54.6%, 32.5–58.8%, and 46.8–66.1% over per-input optimization and 58.4–75.1%, 60.0–78.9%, and 63.2–81.5% over per-workflow optimization, respectively.
  • $\bullet$
    Aragog preserves accuracy compared to always using the largest model, with maximum degradation of 1%.
  • $\bullet$
    Aragog demonstrates robust performance improvements over baselines across diverse model families, router architectures, configuration space sizes, and GPU allocation.
5.1 Experimental Setup
Models. We use the Qwen-2.5 model family [qwen]: 7B, 14B, 32B parameters. This provides a diverse range of capability-cost tradeoffs for each agent in the workflows. We use Snowflake Arctic-Embed-L-v2.0 [snowflake-model] as the router’s embedding model, which has a small number of parameters while providing high-quality embeddings and supporting relatively long context lengths. Additionally, we explore more diverse model family setups and different embedding models in §5.3.
Testbed. We conduct our experiments on an H100 GPU node, which contains 8 NVIDIA H100 80GB GPUs interconnected via NVLink and a 96-core Intel Sapphire Rapids CPU. We use SGLang as our model serving engine. For models larger than 7B parameters, we enable tensor parallelism with a degree of 2 for 14B models and 4 for 32B models, following standard practice. We dedicate the one remained GPU to router inference. For baselines without routing components, this GPU instead serves an additional 7B model to ensure fair resource allocation. We also explore different initial GPU allocation configurations in §5.3. As in prior work [vllm, sglang, autellix], we simulate request arrivals using a Poisson process with varying arrival rates to capture different load conditions.
Workflows and datasets. We evaluate Aragog across diverse workflow patterns and application domains.
  • $\bullet$
    Self-refine for coding on HumanEval [humaneval] (164 programming problems), using a three-agent workflow with generation, critique, and refinement stages.
  • $\bullet$
    Self-refine for QA on SQuAD v2.0 [squad] (1000 sampled questions), following the same three-stage pattern.
  • $\bullet$
    Text-to-SQL on Bird-SQL [bird] (780 samples), where query construction is decomposed into keyword extraction, column selection, SQL generation, and refinement stages.
  • $\bullet$
    Majority voting [self-consistency] (four parallel voters and one aggregator) on MMLU-Pro [mmlu-pro] (1,500 sampled questions), where multiple reasoning paths are generated in parallel and aggregated.
  • $\bullet$
    Task decomposition on StrategyQA [strategyQA] (1,500 sampled questions), where problems first pass through decomposer, then four sub-task solver in parallel, and finally answer aggregator.
Baselines. We compare Aragog against two oracle baselines that represent optimal performance for each category: (1) Per-workflow optimization: selects the cheapest configuration among all configurations that match the most expensive configuration’s accuracy, the configuration stays fixed for all requests; and (2) Per-input optimization: performs router-based selection for each individual request. Both baselines are given an oracle with perfect accuracy knowledge for serving requests, representing the performance upper bound of each approach. Additionally, we enforce the same router inference overhead for the per-input baseline as Aragog. Note that this routing overhead is already underestimated in favor of the per-input baseline’s single monolithic router, as discussed in §3.1. Moreover, all baselines incorporate the same state-of-the-art graph-aware scheduling for agentic workflows as Aragog (§3.2).
Figure 11: Average latency and throughput comparison under varying request rates and across five workflow dataset pairs. Each dot represents the average latency and throughput under a specific request rate.
Figure 12: Median latencies for different workflow dataset pairs under varying request rates, with error bars for P25 and P95 latencies
Metrics. We measure serving capacity (maximal throughput), latency (Average/P25/P50/P90) under varying request rates, and accuracy preservation relative to the most expensive configuration (using the largest model for all agents).
5.2 End-to-End Performance
Figures 1113 compare Aragog with two baselines that represent optimal performance for each category across diverse workloads. Overall, Aragog significantly improves the maximal serving capacity and lowers latencies, while matching the accuracy when always using the largest model.
Serving capacity improvement. In Figure 11, we evaluate the end-to-end serving capacity of Aragog against baselines across five workloads. For each workload, we sweep the incoming request rate from low to high and measure the achieved throughput and average latency at each rate, shown as individual points on the curves. Aragog consistently achieves the highest throughput and lowest average latency under varying request rates. Specifically, Aragog improves maximum serving throughput by 50.0–76.3% over per-input optimization and 78.1–217.0% over per-workflow optimization. Additionally, Aragog reduces average latency under the highest request rates by 25.4–60.0% and 60.0–79.1% over per-input and per-workflow optimization, respectively. These results demonstrate that Aragog scales more effectively under increasing load.
We observe several important trends. First, compared to per-input optimization, Aragog’s capacity improvements grow with workflow complexity. The throughput improvement increases from 50.0–50.1% for self-refine workflows to 54.0–76.3% for more complex workflows with either longer sequential chains (text-to-sql) or parallel branches (task decomposition and voting). This occurs because workflows with more stages exacerbate the limitations of per-input optimization’s static, upfront configuration. The configuration staleness increases with workflow complexity, leading to increasingly suboptimal configuration decisions. In contrast, Aragog exploits runtime reconfiguration and flexible request scheduling to maximize resource utilization across workflow stages, yielding higher throughput gains. Second, compared to per-workflow optimization, Aragog’s capacity improvements are more pronounced for simpler workflows (100.2–217.0%) than for complex workflows (78.1–128.6%). This is because complex workflows offer fewer opportunities for individual requests to exploit cheaper configurations, as tasks are harder. Similarly, the performance gap between per-workflow and per-input optimization narrows as workflow complexity increases.
Latency reduction. Figure 12 presents detailed latency distributions across five workloads under the same request rates used in Figure 11. At low request rates where systems are underutilized, all approaches achieve comparable latency, as expected. However, as load increases, Aragog’s latency reductions across P25, median, and P95 all become increasingly pronounced. Under the highest load across all workloads, Aragog reduces P25, median, and P95 end-to-end latency by 23.6–54.6%, 32.5–58.8%, and 46.8–66.1% over per-input optimization and 58.4–75.1%, 60.0–78.9%, and 63.2–81.5% over per-workflow optimization, respectively. The latency gap widens most dramatically for the P95 metric, where Aragog achieves up to 66.1% and 81.5% reduction over the two baselines. This is because through runtime reconfiguration, Aragog avoids excessive queuing at individual models while maximizing resource utilization, effectively reducing tail latency compared to static baselines.
Accuracy preservation. Figure 13 shows that Aragog maintains accuracy across all five workloads while achieving significant throughput and latency improvements. Aragog’s accuracy matches the most expensive configuration, with differences of at most 1 percentage across all workloads. This demonstrates that Aragog’s configuration predictor maintains high prediction accuracy while providing configuration flexibility for runtime scheduling.
Figure 13: Evaluating Aragog’s end-to-end accuracy against the case when always using the largest model.
5.3 Microbenchmarks
While the end-to-end evaluation demonstrates Aragog’s overall effectiveness, we conduct microbenchmarks to evaluate Aragog under more diverse setups, examining its performance across varying resource allocations, workload characteristics, and system parameters. We present results using the self-refine(HumanEval) and voting(MMLU-Pro) workflows as representative examples; the observed trends hold across all our workloads. Unless otherwise specified, all other experimental settings remain the same as described in §5.1.
System overheads. Aragog’s runtime overhead primarily stems from router inference Figure 14 shows the ratio of average routing latency to total end-to-end latency across varying request rates. Routing overhead remains minimal, accounting for at most 3.5% of total latency at low request rates and decreasing to below 1% at high request rates. Importantly, these percentages reflect the absolute routing latency; much of this overhead is overlapped with each request’s queueing delays. The ratio decreases as the request rate increases because queueing delays grow faster with load compared with routing latency, as routing computation is minimal compared to LLM inference in a workflow. These results demonstrate Aragog’s efficient routing optimizations.
Figure 14: Ratio of average routing latency to total end-to-end latency across workloads and request rates.
Beam search with different beam sizes. Figure 15 shows the impact of beam size on Aragog’s serving capacity improvement over two baselines. For self-refine on HumanEval, throughput improvement over per-input optimization grows from 16.3% to 51.2% as beam size increases from 1 to 8, while improvements over per-workflow optimization increase from 82.2% to 217.1%. Similarly, for voting on MMLU-Pro, improvements rise from 19.2% to 61.7% over per-input and 73.4% to 130.4% over per-workflow optimization. Performance gains plateau with increasing beam size, as beam search finds better configurations than greedy search (beam size 1) but with diminishing returns.
Figure 15: Impact of beam size on serving capacity improvement.
Mixture of Model Family. Figure 16 shows Aragog’s generality to heterogeneous model families by using models from mixed model families: Llama3 8B, Phi4 14B, and Qwen2.5 32B. While achieving the comparable accuracy as the most expensive configuration (within 1% loss), Aragog improves serving throughput by 33.7–68.5% over per-input optimization and 109.5–111.4% over per-workflow optimization. Additionally, Aragog reduces average latency under the highest request rates by 37.9–45.4% and 63.1–63.6% over per-input and per-workflow optimization, respectively. This demonstrates that Aragog generalizes effectively across diverse model architectures.
Figure 16: Average latency and throughput comparison under varying request rates using a mixed model family (Llama 8B, Phi 14B, and Qwen 32B). Each dot represents the average latency and throughput under a request rate.
Table 1: Sensitivity analysis of router’s embedding model, initial resource allocation, and configuration space size
(a) Impact of router architecture
Throughput $\uparrow$ Avg Latency $\downarrow$
Router’s Embedder vs PI vs PW vs PI vs PW
Arctic-Embed-L-v2.0 50.0% 217.0% 60.0% 79.1%
EmbeddingGemma 50.3% 216.8% 60.2% 79.5%
(b) Impact of GPU allocation scheme
Throughput $\uparrow$ Avg Latency $\downarrow$
Resource Allocation vs PI vs PW vs PI vs PW
8 H100 GPUs 50.0% 217.0% 60.0% 79.1%
4 H100 GPUs 28.3% 205.2% 41.2% 78.6%
(c) Impact of configuration space size
Throughput $\uparrow$ Avg Latency $\downarrow$
Configuration Space vs PI vs PW vs PI vs PW
Full Space 50.0% 217.0% 60.0% 79.1%
Constrained Space 34.2% 164.6% 43.2% 51.7%
All improvements measured at the highest request rate from the end-to-end evaluation (§5.2). PI = per-input optimization, PW = per-workflow optimization. Crucially, all experiments maintain the same or better accuracy as the main evaluation.
Impact of router architectures. Table 1(a) compares two embedding models for the router: Arctic-Embed-L-v2.0 (our default) and Google’s EmbeddingGemma [embeddinggemma]. Both models have similar parameter counts and deliver comparable performance. The negligible performance differences demonstrate that Aragog’s routing mechanism is robust to the choice of embedding model, provided the embedding model can capture semantic similarity between workflow requests.
Impact of different GPU allocation schemes. We evaluate Aragog’s robustness to resource constraints by reducing GPU allocation from 8 to 4 H100 GPUs. With fewer GPUs, we reduce tensor parallelism accordingly: Qwen 32B from TP=4 to TP=2, Qwen 14B from TP=2 to TP=1, while Qwen 7B remains on a single GPU. This scenario favors the per-input baseline, which prioritizes smaller models based on static costs, while Aragog’s advantage comes from runtime reconfiguration that shifts load from overloaded small models to underutilized larger models. With 4 GPUs, limited capacity for larger models restricts such reconfiguration opportunities. Table 1(b) shows that Aragog still achieves significant improvements over per-input: 28.3% throughput gain and 41.2% latency reduction with 4 GPUs, compared to 50.0% and 60.0% with 8 GPUs. Gains over per-workflow optimization remain stable (205.2% vs. 217.0% in throughput, 78.6% vs. 79.1% in latency), as it often selects a fixed configuration with large models to meet worst-case accuracy requirements, making it similarly resource-constrained.
Impact of the size of configuration space. Aragog by default explores the full configuration space given a workflow and model candidates. Table 1(c) investigates performance under a constrained space with additional restrictions: for self-refine workflows, the critic and refine agents must use the same model; for voting workflows, all voters must use the same model. With the constrained space, Aragog achieves 34.2% and 43.2% improvements (throughput and latency) over per-input optimization versus 50.0% and 60.0% with the full space, and 164.6% and 51.7% over per-workflow optimization versus 217.0% and 79.1% with the full space. Despite reduced configuration flexibility, Aragog maintains significant advantages through joint scheduling that leverages each request’s available flexibility. This also demonstrates that Aragog’s joint scheduling can accommodate varying levels of configuration flexibility, offering a new dimension (configuration space size) for optimizing larger workflows.
6 Additional Related Work
LLM serving optimizations The rapid deployment of large language models has driven significant innovations in serving optimizations. Some focus on application level compression that craft lightweight variants for efficient inference [awq, gptq, llm-pruner, speculative-decoding]. Some employ compiler techniques to optimize execution for low latency [mirage, mlc-llm]. Other works focus on system perspectives like request scheduling [orca, llmschedulinglearning], KV cache memory management [vllm, jenga], parallelism [alpaserve, nanoflow, magescale-infer], efficient kernels [flashattention, flashinfer], and resource allocation [llumnix, distserve, prism]. Aragog operates as an orchestration layer between LLM serving engines and the agentic workflow layer. As such, these optimizations are orthogonal to Aragog and can be incorporated into the underlying serving engines.
Agent serving optimizations The rise of LLM-powered agents introduces new challenges for LLM serving, as agent execution involves an input-dependent and varying number of interceptions (e.g., tool calls) and LLM calls. Recent research has focused on addressing these agent-specific inefficiencies [autellix, infercept, pdm, tempo, asyncfunccall]. For instance, InferCept [infercept] reduces GPU resource waste arising from frequent LLM interceptions, while Autellix [autellix] mitigates head-of-line blocking caused by the variable number of LLM calls per agent by prioritizing requests based on cumulative service time. These techniques are complementary to Aragog, which orchestrates execution at the workflow level (across a graph of agents) and can leverage agent-serving optimizations within each agent within the graph.
Agentic workflow serving optimizations Agentic workflows have emerged as a more reliable and debuggable approach than fully autonomous agents. These workflows orchestrate multiple agents as a graph, introducing dependencies across agents. Recent systems exploit graph-level knowledge for scheduling and resource management [parrot, ayo, kairos, batchquery]. For instance, Parrot [parrot] enables graph-aware batching and KV cache sharing, while Ayo [ayo] rearranges the execution graph for efficient execution. These systems optimize scheduling and placement for a fixed configuration; Aragog complements them by dynamically adapting the configuration at runtime. Orthogonally, other work automates the design and search of workflow graph structures [aflow, flow]. Aragog is agnostic to the input workflow’s structure and can efficiently serve these automatically generated workflows.
7 Conclusion
We present Aragog, an efficient agentic workflow serving system that progressively adapt request’s configuration to runtime dynamics. Aragog’s core insight is to decouple accuracy and cost based configuration selection. The expensive accuracy-based configuration prediction happens once and before execution, while the lightweight cost-based configuration selection runs frequently throughout execution. Aragog makes runtime reconfiguration efficient with novel acceleration strategies. Across diverse workloads, Aragog achieves 50.0–217.0% higher serving capacity than baselines while maintaining the accuracy of the most expensive configuration. Aragog delivers lower serving latency under different request loads. Under peak load, it delivers 32.5–78.9% and 46.8–81.5% lower median and P95 latency, respectively.